OpenStudy (anonymous):

Solve each equation correct to three decimals. 5^x = 120

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] and a calculator

OpenStudy (anonymous):

Less than 3..

OpenStudy (anonymous):

Thank you very much for the equation, I understand the problem now.

OpenStudy (anonymous):

ln(5^x) = ln(120) xln5 = ln150 x = ln(120)/ln(5) Put that in a calculator and see what it gives you, I get 1.609